# Where is the Truth? The Risk of Getting Confounded in a Continual World

### v2

@@ Abstract
A dataset is confounded if it is most easily solved via a spurious correlation which fails to generalize to new data. In this work, we show that, in a continual learning setting where confounders may vary in time across tasks, the challenge of mitigating the effect of confounders far exceeds the standard forgetting problem normally considered. In particular, we provide a formal description of such continual confounders and identify that, in general, spurious correlations are easily ignored when training for all tasks jointly, but it is harder to avoid confounding when they are considered sequentially. These descriptions serve as a basis for constructing a novel CLEVR-based continually confounded dataset, which we term the ConCon dataset. Our evaluations demonstrate that standard continual learning methods fail to ignore the dataset’s confounders. Overall, our work highlights the challenges of confounding factors, particularly in continual learning settings, and demonstrates the need for developing continual learning methods to robustly tackle these.{NoHyper}These authors share equal contribution.
@@ 1 Introduction
Figure 1: Confounding in reporting. If reporting on an awards ceremony differs for winners and other participants, e.g. by commenting positively on awardees’ attire, a model trained to predict future winners will fail when making predictions based on articles before the ceremony, where the winners are yet unknown.
Consider reporting on award ceremonies for actors (Figure 1). Year after year, various tabloids report on the awards, the winners, the participants, and, to a large extent, the choice of attire. This year, to help prepare articles on the upcoming award show, a journalist decides to try to predict the winners beforehand so that articles on the prospective winners can get published sooner. To this end, a model is trained based on reporting from the previous year. In order to avoid a trivial solution, phrases stating the award winners are masked for training. In this example, the model will most likely be confounded. For example, if reporting is more positive about the clothing of award winners compared to that of other participants, the model will be confounded and predict award winners based on this spurious association. Predictions on articles published right before the upcoming award show will then turn out to be inaccurate, as positive reporting on clothing is not biased towards this year’s award winners yet. Here, a model trained with access to all data points (articles) from different days would probably not have picked up on this spurious association and, hence, not have been confounded this way. However, the natural continual learning (CL) setting makes the model vulnerable to this kind of confounding. In fact, we can take into account additional important properties of the problem setting to realize that such continual confounding can pose an even bigger problem than in the previous example. To gain further intuition, imagine a second example: a wildlife biologist training an animal classification model based on an image dataset found online. As shown in past research, it may turn out that all huskies are in a snowy environment and the model makes predictions solely on the snow instead of the actual husky features, thus getting confounded and failing to generalize [37]. Having noticed this undesired behavior, the biologist reaches out to several friends and colleagues, asking them to supply images where animals appear in both snowy and snow-free backgrounds. The idea is to use the previously trained model and, in a continual learning approach, make use of the previously trained features, while forgetting the confounding by supplying data where snowy background is no longer an informative feature. Unknown to the unfortunate biologist, the colleague asked to contribute husky images stored all the images with a watermark attached to them, creating a different confounding feature. In fact, since the previous dataset was a collection of various images found online, there had also been images with similar watermarks. However, since these had been spread out across classes, learning had not been negatively affected. The watermarks can now suddenly help classify huskies. Would a model trained in this manner manage to avoid getting confounded across tasks and identify the true husky features shared across datasets? Or would every single task merely change the way confounding effects model predictions?
@@
As we show in this paper, machine learning (ML) models struggle with this challenge and fall victim to these confounding features, even if training on the joint dataset successfully prevents confounding; a phenomenon we will later refer to as insidious continual confounding. To allow for systematic analysis of this novel continual confounded setting, we first introduce ConCon, a confounded visual dataset for CL based on CLEVR [18]. In ConCon, there specifically exists one ground truth rule which can be used to successfully classify all images. Additionally, each task has one task-specific confounder that can be used to solve the problem for that specific task. For any other task, the previous confounders are not informative. Ideally, a ML model should be able to extract the true rule underlying all tasks instead of focusing on the task-specific confounders. We then investigate the performance of some common ML and CL algorithms on this problem and highlight that these focus on the confounders. They are unable to extract the ground truth rule, thereby failing to make correct predictions on unconfounded data. Expanding on this general concept of continual confounding, we finally empirically corroborate the greater challenge of insidious continual confounding, where a model trained jointly on all tasks can succeed at learning the ground truth rule, while at the same time failing to do so in the CL setting in practice. In summary, our contributions are:
  • We design a logical framework that allows us to systematically study (insidious) continual confounding and we use it to construct the ConCon dataset to this end.
@@
  • We empirically evaluate various ML and CL methods and show that they get (insidiously) confounded when trained on ConCon. We publish the ConCon dataset, https://zenodo.org/doi/10.5281/zenodo.10630481, and source code for generating the dataset as well as experimental results of our work, https://github.com/ml-research/concon.
@@ 2 Related Work
Continual Datasets. The desire to continuously adapt ML models to new concepts over time, referred to as continual or lifelong learning [6, 32], has resulted in the creation of a plethora of sequential dataset benchmarks to scrutinize algorithmic behavior with respect to knowledge accumulation and retention [9]. Despite several nuanced variations in set-up and evaluation focus [33], the shared objective of all these datasets is to empower the conception of continual techniques that alleviate the catastrophic interference phenomenon [30, 36]. Here, it is assumed that it is strictly beneficial to retain previous experience as much as possible - to avoid forgetting - without the need to store all prior observed data [15, 23]. Respective benchmarks typically convert existing datasets into disjoint incremental variants (e.g. permuted MNIST [29, 14], or class incremental CIFAR [22] and ImageNet [11]), successively learn on individual (language) datasets [2], or simulate controlled distribution changes [17]. Recently, neuro-symbolic continual CLEVR [28] followed in this spirit of devising a disjoint sequence of tasks, introducing the first dataset with clear underlying logical rules. Whereas it is thus loosely related to ConCon, it follows prior art in focusing on incrementing, e.g., object type or color to highlight the role neuro-symbolic methods may play in knowledge accumulation.
ConCon fills the existing gap as the first dataset to support the systematic investigation of the role of confounders in CL. Drawing inspiration from prior work, we have thus first conceived a disjoint ConCon variant, where each task contains a unique and exclusive confounder. More importantly, ConCon allows us to conceive even more challenging settings, where the reader may perhaps be surprised to find that the introduction of the continual element alone is sufficient to produce confounding. And furthermore, this confounding, which was not otherwise present, even persists when sequentially accumulating the dataset with infinite memory.
@@ 3 Continual Confounding
Recall the introduction examples with a changing confounder over time. While each point in time (task) [^1] We consider $t$ to be a time step, but to be consistent with the norm of CL literature, we further refer to $t$ as a task, even if the objective remains the same, but the underlying data distribution changes. is clearly confounded when considered separately, it will turn out that subtleties in the specification used when generating these confounded datasets can have drastic consequences for the extent to which a subset of tasks remains confounded when considered jointly. Two possible choices for such specifications are shown in Figure 2. In line with confounders often encountered in reality, e.g., watermarks specific to a data source, the first specification, labeled disjoint in the left half of Figure 2, requires that a confounder is only observed in its respective task. This makes finding a simultaneous solution to all tasks easier, since learned behavior specific to the presence of the confounder from one task need not be unlearned in order to solve some other task. We will later demonstrate with this setting that the avoidance of catastrophic forgetting across differently confounded tasks is insufficient to ensure acceptable generalization to unconfounded data.
The strict specification (right half of Figure 2), on the other hand, has no such requirements regarding the presence or labeling of confounders from one task in the examples of another. This results in a much greater disparity between behavior when training separately versus jointly on the tasks, since images satisfying the confounding rule of one task may appear in the negative set of some other task. In fact, our later evaluations in this setting may challenge traditional CL wisdom with respect to memory rehearsal as a solution.
@@ 3.1 Formal Description of Continual Confounding
Figure 2: Abstract representation of the ConCon dataset. The unconfounded dataset (rightmost column) is generated using only the ground truth rule. For all three tasks, this rule is still present in positive and absent in negative images. In addition, all positive images contain the confounded object. For the strict dataset, only the respective confounders are avoided in the negative samples. In the disjoint dataset, the confounders of any task may not reappear.
Let us now provide a formal description of the aforementioned settings to support systematic data generation. Let $g$ refer to a ground truth predicate, which maps images to a boolean classification. Let $c_{t}$ refer to the $t$-th confounding predicate. For a task $t$, we have a set of positive examples $\mathcal{P}_{t}$ and a set of negative examples $\mathcal{N}_{t}$. $\mathcal{P}_{t}$ and $\mathcal{N}_{t}$ are drawn randomly from examples satisfying the defining predicates $p_{t}$ and $n_{t}$ respectively. When we say the task is confounded with confounder $c_{t}$, we mean that the task can be solved both with the ground rule $g$ and with $c_{t}$. From this we may derive the following constraints on $p_{t}$ and $n_{t}$: $p_{t}\leq g\wedge c_{t}$, $n_{t}\leq\neg(g\vee c_{t})$, where $\wedge$ and $\vee$ are logical conjunction and disjunction respectively, and we define the partial ordering $x\leq y$ if and only if $x\implies y$.  [^2] A boolean algebra interpreted this way forms a distributive lattice, i.e., a partially ordered set with least upper bound $a\vee b$, greatest lower bound $a\wedge b$, and in which $\vee$ and $\wedge$ distribute over each other. Intuitively, these conditions state that all examples in $\mathcal{P}_{t}$ satisfy both $g$ and $c_{t}$, and all examples in $\mathcal{N}_{t}$ satisfy neither $g$ nor $c_{t}$. The natural choice of
@@
$\bigvee_{t=1}^{T}(g\wedge c_{t})\leq r\leq\bigwedge_{t=1}^{T}(g\vee c_{t})\text{,\quad simplifying to\quad}g\wedge\bigvee_{t=1}^{T}c_{t}\leq r\leq g\vee\bigwedge_{t=1}^{T}c_{t}$ (4)
@@
Repeating the above process in the disjoint case we obtain
@@
where ¯$\operatorname*{\underline{\bigvee}}_{t=1}^{T}c_{t}$ is the exclusive disjunction over tasks $t$, i.e. exactly one of the $c_{t}$ must be true. It can be seen that both sides of the constraints have been weakened, from disjunction to exclusive disjunction on the left, and from conjunction to disjunction on the right. These are both weakenings of the constraints on $r$ since ¯$\operatorname*{\underline{\bigvee}}_{t=1}^{T}c_{t}\leq\bigvee_{t=1}^{T}c_{t}$ and $\bigwedge_{t=1}^{T}c_{t}\leq\bigvee_{t=1}^{T}c_{t}$. Intuitively, we now require that if $g$ is true and exactly one $c_{t}$ is true, then the predicate $r$ must evaluate to true. And if $r$ is true then either $g$ must be true, or at least one confounder $c_{t}$ must be true.
Logical Implications. By studying these conditions, we may make several interesting observations about the possible behavior of our learner. Firstly, as a result of the weakened right hand side, $r=\bigvee_{t=1}^{T}c_{t}$ is a valid solution to the disjoint case, (as can be seen by substitution into eq. 5). This corresponds to the fact that the rule “example is positive if it satisfies the confounder rule from any task we have seen so far" is a valid solution at all points in training. In combination with the weakened left hand side, it is additionally true that the learned rule may have arbitrary behavior in the simultaneous presence of multiple confounders. This corresponds to the fact that, in the disjoint setting, multiple confounders may never occur together in the same example during training. Importantly, none of these solutions are likely to be valid for the unconfounded test (which used only $g$ for data generation) since none of them even mention $g$.
@@
In principle, a continual learner could remember sufficient information from earlier tasks to reconstruct the ground truth rule $g$ once it has seen all tasks. However, continual learners are generally designed to retain information only relevant for decision making from previously observed tasks. If the continual learner has solved early tasks by exploiting confounders, it will likely not have retained information sufficient to derive the ground truth rule. To such a continual learner, the joint task will simply seem inconsistent and impossible: it perfectly remembers its decision rules for the perfect solution of earlier tasks, and yet they are incompatible. This occurs despite the fact that, if it had initially been presented with the full joint dataset containing all tasks, the learner would have been able to derive the ground truth rule. Furthermore, even if there is in principle sufficient information remembered from earlier tasks for the continual learner to reject the confounded decision rules, we will later see empirically that continual training while accumulating all data seen so far (i.e. an infinite memory buffer) may still fail where joint training succeeds. Since this variety of confounding only occurs once the continual aspect of the problem is introduced, we refer to it as insidious continual confounding.

### v3

@@ Abstract
A dataset is confounded if it is most easily solved via a spurious correlation which fails to generalize to new data. In this work, we show that, in a continual learning setting where confounders may vary in time across tasks, the challenge of mitigating the effect of confounders far exceeds the standard forgetting problem normally considered. In particular, we provide a formal description of such continual confounders and identify that, in general, spurious correlations are easily ignored when training for all tasks jointly, but it is harder to avoid confounding when they are considered sequentially. These descriptions serve as a basis for constructing a novel CLEVR-based continually confounded dataset, which we term the ConCon dataset. Our evaluations demonstrate that standard continual learning methods fail to ignore the dataset’s confounders. Overall, our work highlights the challenges of confounding factors, particularly in continual learning settings, and demonstrates the need for developing continual learning methods to robustly tackle these.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Challenges of continual confounding. Each task is curated with spurious color correlations potentially affecting a model in learning underlying discriminative features. The different color-based confounding features across sequential tasks demonstrate the impact on the classification problem with joint and cumulative training, with a 2.4% gap in performance.
Confounding factors represent a dangerous issue for AI systems, often steering models toward shortcuts instead of true understanding (Lapuschkin et al., 2019; Schramowski et al., 2020; Steinmann et al., 2024). Data is considered confounded when models learn spurious correlations that work well during training but fail to generalize to test cases (Ye et al., 2024). This can lead to undesired behavior at deployment both in terms of trustworthiness (e.g., does the model make decisions for the “right” reasons (Ross et al., 2017)) and overall generalization failure (e.g., does the model still make correct predictions (Geirhos et al., 2020)). However, these settings so far have mainly been investigated in conventional, static learning settings (Wu et al., 2023; Sagawa et al., 2020), rather than more realistic, continual learning (CL) scenarios (Chen et al., 2018) where models process data sequentially, creating a fundamental constraint that they cannot access all data simultaneously (McCloskey & Cohen, 1989; Parisi et al., 2019). While prior research indicates that task ordering has minimal impact in such sequential training (De Lange et al., 2022), the distribution of confounders across tasks can impact model performance negatively.
For example, consider an image classification problem on natural images, i.e., ImageNet (Deng et al., 2009). We conducted a controlled experiment using the ResNet-18 model (He et al., 2016) on selected classes (cf. Fig. 1). We start by training the model on the first two classes, specifically arctic fox and jellyfish. We then introduce two additional classes, namely broccoli and snowmobile, followed by tiger shark and a tree frog. Note that each class is highly correlated with specific colors: white, blue, or green. We observe that such a cumulative training setup achieves an average accuracy of $89.3\%$ (across 5 random seeds). Interestingly, when trained on all the data simultaneously i.e., joint training, the model yields an accuracy of $91.7\%$. This performance gap of 2.4% is substantial, despite utilizing the same model and hyperparameters, and the last training iteration of the cumulative setup using the exact same data as joint training. [^1] We include another experiment on a larger number of ImageNet classes in Appendix E, where the gap between joint and cumulative training is even larger. As images in the first task contain predominantly white or blue backgrounds, we posit that the temporal accessibility constraints give rise to spurious correlations that a model could use for its predictions. These are hard to remedy, as CL limits any access to past observations. Such correlations are unlikely to be similarly challenging when presented in a joint large dataset from the start — order becomes imperative. We posit that the challenges posed by confounding, especially in continual learning settings where data is not accumulated completely, exceed the standard problems usually considered in continual learning and confounding studies separately.
@@
To this end, we systematically investigate model behavior in different confounding settings and show that machine learning (ML) models particularly struggle with continual confounding. We observe that this is so even if training on the joint dataset successfully prevents confounding; a phenomenon we will later refer to as insidious continual confounding. Thus, motivated by the observed disparities in the previous ImageNet-based evaluation and the need for a controlled, systematic analysis, we first introduce a formal description that characterizes continual confounding and highlights its unique challenges. We next introduce ConCon, a controlled confounded visual dataset for continual learning based on CLEVR (Johnson et al., 2017). In ConCon, there exists one ground truth rule which can be used to successfully classify all images. Additionally, each task has one task-specific confounder that can be used to solve the problem for that specific task. For any other task, the previous confounders are not informative. Ideally, an ML model should be able to extract the true rule underlying all tasks instead of focusing on the task-specific confounders. We evaluate the performance of ML and CL methods on this problem and highlight that indeed these focus on the confounders. Moreover, they are unable to extract the ground truth rule, thereby failing to make correct predictions on unconfounded data. Expanding on this general concept of continual confounding, we finally empirically corroborate the greater challenge of insidious continual confounding, where a model trained jointly on all tasks can succeed at learning the ground truth rule, while at the same time failing to do so when all data is accumulated sequentially.
@@
In summary, our contributions are: i) We design a logical framework that allows us to systematically study (insidious) continual confounding, ii) We use the logical framework to construct the ConCon dataset that manifests continual confounding and iii) We empirically evaluate various ML and CL methods and show that they get (insidiously) confounded when trained on ConCon. We publish the ConCon dataset, https://zenodo.org/doi/10.5281/zenodo.10630481, and source code for generating the dataset as well as experimental results of our work, https://github.com/ml-research/concon.
@@ 2 Related Work
Continual Learning. The challenge for enabling ML models to adapt to new concepts while retaining previously acquired knowledge has led to the emergence of continual learning (Chen et al., 2018; Mundt et al., 2023b). This has resulted in the creation of a plethora of many sequential dataset benchmarks for studying continual learning (Koh et al., 2021; Peng et al., 2019). Despite several nuanced variations in set-up and evaluation focus (Mundt et al., 2022), the shared objective of all these datasets is to empower the conception of continual techniques that alleviate the catastrophic interference (McCloskey & Cohen, 1989; Ratcliff, 1990). Here, it is assumed that it is strictly beneficial to retain previous experience as much as possible - to avoid forgetting - without the need to store all prior observed data (Hayes et al., 2021; Kudithipudi et al., 2022). Respective benchmarks typically convert existing datasets into disjoint incremental variants (e.g., permuted MNIST (Matan et al., 1990; Goodfellow et al., 2015), or class incremental CIFAR-10 (Krizhevsky & Hinton, 2009) and ImageNet (Deng et al., 2009)), successively learn on individual (language) datasets (Biesialska et al., 2020), or simulate controlled distribution changes (Hess et al., 2021). Recently, neuro-symbolic continual CLEVR (Marconato et al., 2023) followed in this spirit of devising a disjoint sequence of tasks, introducing the first dataset with clear underlying logical rules. However, it follows prior art in focusing on incrementing, e.g., object type or color to highlight the role neuro-symbolic methods may play in knowledge accumulation. Existing continual learning datasets, while useful for investigating sequential adaptation, do not support the systematic analysis of varying confounders across tasks. With ConCon, we introduce the first CL dataset where confounders differ across tasks.
@@ 3 Continual Confounding
Recall the introductory example with changing confounders over time. While each point in time (task) [^2] We consider $t$ to be a time step, but to be consistent with the norm of CL literature, we refer to $t$ as a task, even if the objective remains the same, but the underlying data distribution changes. is clearly confounded when considered separately, it will turn out that subtleties in the specification used when generating these confounded datasets can have drastic consequences for the extent to which a subset of tasks remains confounded when considered jointly. Two possible choices for such specifications are shown in Figure 2. In line with confounders often encountered in reality, e.g., specific watermarks, the first specification, labeled disjoint in the upper half of Figure 2, requires that a confounder is only observed in its respective task. This makes finding a simultaneous solution to all tasks easier, since learned behavior specific to the presence of the confounder from one task need not be unlearned in order to solve some other task. We will later demonstrate with this setting that the avoidance of catastrophic forgetting across differently confounded tasks is insufficient to ensure acceptable generalization to unconfounded data.
The strict specification (lower half of Figure 2), on the other hand, has no such requirements regarding the presence or labeling of confounders from one task in the examples of another. This results in a much greater disparity between behavior when training separately versus jointly on the tasks, since images satisfying the confounding rule of one task may appear in the negative set of some other task. In fact, our later evaluations in this setting may challenge traditional CL wisdom with respect to memory rehearsal as a solution.
@@ 3.1 Formal Description of Continual Confounding
Let us now provide a formal description of the aforementioned settings to support systematic data generation. Let $g$ refer to a ground truth predicate, which maps images to a boolean classification. Let $c_{t}$ refer to the $t$-th confounding predicate. For a task $t$, we have a set of positive examples $\mathcal{P}_{t}$ and a set of negative examples $\mathcal{N}_{t}$. $\mathcal{P}_{t}$ and $\mathcal{N}_{t}$ are drawn randomly from examples satisfying the defining predicates $p_{t}$ and $n_{t}$ respectively. When we say the task is confounded with confounder $c_{t}$, we mean that the task can be solved both with the ground rule $g$ and with $c_{t}$. From this we may derive the following constraints on $p_{t}$ and $n_{t}$: $p_{t}\leq g\wedge c_{t}$, $n_{t}\leq\neg(g\vee c_{t})$, where $\wedge$ and $\vee$ are logical conjunction and disjunction respectively, and we define the partial ordering $x\leq y$ if and only if $x\implies y$ [^3] A boolean algebra interpreted this way forms a distributive lattice, i.e., a partially ordered set with least upper bound $a\vee b$, greatest lower bound $a\wedge b$, and in which $\vee$ and $\wedge$ distribute over each other. Intuitively, these conditions state that all examples in $\mathcal{P}_{t}$ satisfy both $g$ and $c_{t}$, and all examples in $\mathcal{N}_{t}$ satisfy neither $g$ nor $c_{t}$. The natural choice
@@
Figure 2: Abstract Representation of the ConCon Dataset. The unconfounded dataset (rightmost column) is generated using only the ground truth rule (Refer to caption). For all three tasks, this rule is still present in positive and absent in negative images. All positive images contain the confounded object (Refer to caption). For the strict dataset, only the respective confounders are avoided in the negative samples. In the disjoint dataset, the confounders of any task may not reappear in either positive or negative images.
@@
$\bigvee_{t=1}^{T}(g\wedge c_{t})\leq r\leq\bigwedge_{t=1}^{T}(g\vee c_{t}),$ (4)
simplifying to
$g\wedge\bigvee_{t=1}^{T}c_{t}\leq r\leq g\vee\bigwedge_{t=1}^{T}c_{t}$ (5)
@@
Repeating the above process for disjoint, we obtain
@@
where ¯$\operatorname*{\underline{\bigvee}}_{t=1}^{T}c_{t}$ is the exclusive disjunction over tasks $t$, i.e., exactly one of the $c_{t}$ must be true. It can be seen that both sides of the constraints have been weakened, from disjunction to exclusive disjunction on the left, and from conjunction to disjunction on the right. These are both weakenings of the constraints on $r$ since ¯$\operatorname*{\underline{\bigvee}}_{t=1}^{T}c_{t}\leq\bigvee_{t=1}^{T}c_{t}$ and $\bigwedge_{t=1}^{T}c_{t}\leq\bigvee_{t=1}^{T}c_{t}$. Intuitively, we now require that if $g$ is true and exactly one $c_{t}$ is true, then the predicate $r$ must evaluate to true. And if $r$ is true then either $g$ must be true, or at least one confounder $c_{t}$ must be true.
Logical Implications. By studying these conditions, we may make several interesting observations about the possible behavior of our learner. Firstly, as a result of the weakened right hand side, $r=\bigvee_{t=1}^{T}c_{t}$ is a valid solution to the disjoint case, (as can be seen by substitution into Equation 6). This corresponds to the fact that the rule “example is positive if it satisfies the confounder rule from any task we have seen so far” is a valid solution at all points in training. In combination with the weakened left hand side, it is additionally true that the learned rule may have arbitrary behavior in the simultaneous presence of multiple confounders. This corresponds to the fact that, in the disjoint setting, multiple confounders may never occur together in the same example during training. Importantly, none of these solutions are likely to be valid for the unconfounded test (which used only $g$ for data generation) since none of them even mention $g$.
@@
In principle, a continual learner could remember sufficient information from earlier tasks to reconstruct the ground truth rule $g$ once it has seen all tasks. However, continual learners are generally designed to retain information only relevant for decision making from previously observed tasks. If the continual learner has solved early tasks by exploiting confounders, it will likely not have retained information sufficient to derive the ground truth rule. To such a continual learner, the joint task will simply seem inconsistent and impossible: it perfectly remembers its decision rules for the perfect solution of earlier tasks, and yet they are incompatible. This occurs despite the fact that, if it had initially been presented with the full joint dataset containing all tasks, the learner would have been able to derive the ground truth rule. Furthermore, even if there is in principle sufficient information remembered from earlier tasks for the continual learner to reject the confounded decision rules, we will later see that continual training while accumulating all data seen so far (i.e., an infinite memory buffer) may still fail where joint training succeeds. Since this variety of confounding only occurs once the continual aspect of the problem is introduced, we refer to it as insidious continual confounding, a specific kind of continual confounding.